# Elastohydrodynamics of 3D chemically active filaments

### v1

@@ Abstract
Active deformable filaments exhibit a large range of qualitatively different three-dimensional dynamics, depending on their flexibility, the strength of the active forcing, and the surrounding environment. For a sufficiently strong activity, the resulting forces buckle the filament into different shapes, altering its dynamic behaviour. We investigate the dynamic behaviour of elastic, chemically propelled phoretic filaments, combining two existing models; a local version of slender phoretic theory, which determines the resulting slip flows for chemically propelled filaments with a given shape and chemical patterning, is paired with an efficient method for capturing the elastohydrodynamics of a deformable filament in viscous flow to study the chemoelastohydrodynamics of filaments. As the activity increases, or equivalently the filament stiffness decreases, these filaments undergo buckling instabilities that alter their behaviour from rigid rods. We follow their behaviour well beyond the buckling threshold to find a rich array of dynamics, including motions that may be described as planar swimming, circling, pinwheeling, flapping and diffusive.
@@ 1 Introduction
Active colloids have a prominent position as self-propelling agents that enable the study of multi-agent systems, called ‘active matter’ (Ramaswamy, 2010). These canonical systems of active matter serve as micro-engineered analogues of crowds, flocks of birds or bacterial swarms. A common propulsion mechanism is phoresis, where particles move in gradients of a surrounding field (Anderson, 1989). This occurs due to short-range interactions between the particle’s surface and the surrounding field that generate an effective slip flow close to the particle’s surface, such as solute concentration (Golestanian et al., 2005), temperature (Bickel et al., 2013) or electric field (Nourhani et al., 2015). If the particle also creates gradients in the surrounding propulsive field, then it can self-propel, called autophoresis (Paxton et al., 2004).
Chemically active (diffusiophoretic) colloids are microscale particles that swim by converting a solute ‘fuel’ in their environment into propulsive slip flows (Golestanian et al., 2007). The canonical chemically active colloid is the ‘Janus’ particle, which is half active and half inert; typically a rigid sphere, rod, or disk, partially coated in a catalyst, such as platinum, in a surrounding solute such as hydrogen peroxide solution (Howse et al., 2007; Ebbens & Howse, 2011). Differential reaction rates at the particle’s surface create surface gradients in the solute concentration; the resulting short-range solute-surface interactions generate pressure gradients in a thin layer that drives slip flows close to the surface that propel the swimmer.
@@
Microscale flexible filaments in viscous flows are well-studied, due in part to their prevalence in many biological and artificial systems. In biology, carpets of slender filaments known as cilia actively pump fluid over cycles driven by internal active forces; sperm cells swim due to active waves sent down their long and thin tail-like appendages; and bacteria rotate slender, helical flagella to propel themselves (Lauga & Powers, 2009; Katsamba & Lauga, 2019). Meanwhile, flexible fibres are present in many industrial and household processes, such as textile fibres shed during clothes washing, wood pulp fibres in the production of paper, and microplastics that can pollute the environment (Lundell et al., 2011).
The dynamics of these flexible filaments in viscous flows are difficult to simulate since there is intrinsic feedback between the filament shape, the distribution of forces and the surrounding fluid. There are also accompanying constraints, such as inextensiblity, that must be satisfied throughout the evolution of the system. On top of this, simulations are often computationally expensive because the underlying equations are notoriously numerically stiff (Du Roure et al., 2019). A range of methods have been used to model and simulate the hydrodynamics of slender flexible filaments. Local slender body theories, such as resistive force theory, (Hancock, 1953; Gray & Hancock, 1955; Cox, 1970; Batchelor, 1970; Lighthill, 1976) have been applied to generate fast, efficient simulations of elastic filaments (Moreau et al., 2018; Walker et al., 2020), while non-local slender body theories (Keller & Rubinow, 1976; Johnson, 1977, 1980; Koens & Lauga, 2018) provide greater accuracy through more intensive simulations (Tornberg & Shelley, 2004). Alternative approaches involve using singularity methods such as regularised Stokeslets (Cortez, 2001; Cortez et al., 2005; Hall-McNair et al., 2019), immersed boundary methods (Pozrikidis, 1992; Peskin, 2002) and bead models (Delmotte et al. (2015) and references therein). These models have been developed further to investigate the interaction of flexible fibres and the behaviour of passive and active suspensions (Tornberg & Shelley, 2004; Schoeller et al., 2021).
@@
Active fibres have also been found to undergo similar dynamic transitions depending on the strength of activity relative to the fibre stiffness, with a particular focus on clamped filaments as models of ciliary beating or actin filaments. Bead-spring models of a chain of elastically linked beads, such as with stresslets distributed along its axis (Laskar et al., 2013) or when forced at one end (Laskar & Adhikari, 2017), demonstrated whirling, corkscrewing and planar beating dynamics; similar linear, helical and planar modes were found when attached to a passive cargo (Manna et al., 2017). This bead-spring approach has recently been extended to finer resolution to investigate active poroelastic filaments under follower forces (Altunkeyik et al., 2025), which showed rich transitions between dynamic behaviours, including for chemically active poroelastic filaments, which exhibited regimes of periodic and diffusive-like motion.
@@
An alternative approach to studying active elastohydrodynamics is to combine resistive force theory with the mechanics of elastic filaments. These models have been used to show that clamped filaments under follower forces undergo Hopf bifurcations, seen as a flapping motion (De Canio et al., 2017), while a two-dimensional linear stability analysis of this model revealed transitions to rotation or beating that depends on mechanical boundary conditions, applied load, and noise effects (Fily et al., 2020). Further dynamical regimes have been explored using this elastohydrodynamic theory when applying localized tangential forces (Man & Kanso, 2019) to extract motions such as straight swimming, circling, spiralling and undulating, and also with clamped filaments in confined domains (Stein et al., 2021), revealing transitions of filaments with distributed tangential forces from stable to oscillatory and streaming modes. Recent theoretical contributions by Clarke et al. (2024) and Schnitzer (2025) further investigate the follower-force models, identifying beating and whirling onset via bifurcation theory and weakly nonlinear analysis, enriching the fundamental understanding of these dynamics. In a related system, Lough et al. (2023) modelled the swarmer cell state of Proteus mirabilis as an active Kirchoff rod, and found different swimming modes that may be planar or twisted into three-dimensions, depending on the relative material stiffnesses. Although many of these models abstract away the details provided by the surface chemistry of phoretic propulsion, for example by using follower forces, we expect that they retain essential features of force-induced shape transformations, complex locomotion, and highlight the complex interplay of elasticity, active forcing, and boundary conditions in shaping filament dynamics.
@@
In this paper, we combine recent advances in the efficient simulation of elastohydrodynamic filaments (Walker et al., 2020) with the asymptotically-derived framework for calculating slip flows in chemically active, diffusiophoretic filaments (Katsamba et al., 2020, 2022, 2024) to investigate the fundamental dynamic behaviours of individual chemically propelled elastic filaments. We consider examples of canonical surface chemical patterning to study the transition between qualitatively different dynamic regimes as the material stiffness is varied. Section 2 outlines our theoretical method and numerical implementation for modelling chemically active filaments, using a local slender body theory to calculate the resulting slip flows from chemical reactions on the filament’s surface, which are applied as a background flow in the numerical simulations of the elastohydrodynamics. We present our results in section 3, investigating typical behaviours as the material stiffness (or equivalently the activity) is varied, via two key examples: a symmetrically-patterned filament that buckles under its own induced slip flow, and an asymmetrically-patterned filament whose swimming trajectory is altered by its deformation. We discuss and analyse the outcomes further in section 4.
@@ 2 Theory of chemoelastohydrodynamic filaments
We consider an inextensible slender elastic filament in an otherwise quiescent, infinite viscous fluid, as shown in fig. 1. Within the fluid is a chemical solute that diffuses rapidly. This solute acts as a fuel for the filament’s self-propulsion: a catalytic coating on the filament surface promotes a reaction that locally alters the solute, and the short-range interactions between solute molecules and the filament’s surface generate propulsive slip flows from solute concentration gradients within a thin layer by the filament’s surface.
This phoretic slip flow induces bulk flow in the surrounding fluid, which can result in filament propulsion, but this flow also applies viscous forces on the filament that can deform it. Any change in the filament shape can alter the generated slip flow and thereby affect its motion; there is a complex feedback between shape, chemical patterning and swimming trajectory that we wish to explore.
Here, we outline the main theoretical components used to capture the evolution of a chemically propelled elastic filament. Our work combines two pre-existing methods: Slender Phoretic Theory (Katsamba et al., 2020, 2022, 2024; Katsamba & Montenegro-Johnson, 2024) that asymptotically calculates the leading-order slip flow around slender chemically active filaments when given their chemical patterning and shape, and an efficient method for simulating elastohydrodynamics (Walker et al., 2020). Together, these allow us to probe the dynamic behaviour of chemically propelled elastic filaments, which we refer to as chemoelastohydrodynamics. The fundamentals of these methods are given below.
@@ 2.1 Local slender phoretic theory
We consider a slender filament with a circular cross-section, whose geometry is captured by its centreline position, $\bm{X}(s,t)$, and cross-sectional radius $\rho(s)$, each parametrised by the arclength $s$ such that $|\bm{X}_{s}|=1$. The filament centreline has a total length $2l$ and the maximum cross-sectional radius is $r_{f}$, and we assume the filament is unshearable, so that cross-sections remains perpendicular to the centreline. We define a slenderness parameter by the ratio of the maximum diameter to the total arclength, $\epsilon=r_{f}/l$, which can be considered an aspect ratio, and we assume that our filaments are long and thin so that $\epsilon\ll 1$.
A chemical reaction of the solute dispersed in the surrounding fluid is catalysed by the catalytic coating on the surface of the filament. The resulting solute concentration gradients drive surface slip flows and hence the self-propulsion of the chemoelastic filament. Consequently, there are two important surface chemical properties of the filament: the activity, $\mathcal{A}(\bm{x})$, which prescribes a surface solute flux (either generating or depleting solute), and the mobility, $\mathcal{M}(\bm{x})$, that quantifies how chemical gradients drive to slip flows. Theoretically, we can set these independently and they may vary over the surface.
The system is non-dimensionalised: lengths along the filament are rescaled by the half-length of the filament $l$, while cross-sectional distances are rescaled by the maximum radius $r_{f}$. Taking typical (maximum) values for the activity and mobility as $A$ and $M$, respectively, we also use a concentration scale $Ar_{f}/D$ and a velocity scale $AMr_{f}/Dl$ to rescale our equations, where $D$ is the diffusivity of the solute. All equations are henceforth given in dimensionless form.
Full details of Slender Phoretic Theory are given in Katsamba et al. (2020, 2022, 2024); Katsamba & Montenegro-Johnson (2024). We consider the zero Péclet number limit for the solute dynamics, where diffusion dominates advection, so that the solute concentration, $c$, is governed by Laplace’s equation,
@@
where we consider the disturbance from a background concentration, so that $c\to 0$ far from the surface. The activity provides a normal flux boundary condition at the filament surface via zeroth order reaction kinetics, and any resulting surface concentration gradients drive the slip flow proportional to the mobility. These are boundary conditions on the surface of the filament
@@
where $\bm{n}_{f}$ is the outward normal to the solid surface. Importantly, we frame our results in terms of the arclength $s$ and local polar coordinates in the cross-section $(r,\theta)$. Surface properties are therefore defined by the arclength $s\in[-1,1]$ along the centreline and angle $\theta\in[-\pi,\pi]$ around the cross-section, e.g. $\mathcal{A}=\mathcal{A}(s,\theta)$ and $\mathcal{M}=\mathcal{M}(s,\theta)$, with the surface at $r=\rho(s)\in[0,1]$. These local coordinates are shown in the inset to fig. 1.
In the framework of Slender Phoretic Theory (SPT) (Katsamba et al., 2020, 2022, 2024; Katsamba & Montenegro-Johnson, 2024), the solution of Laplace’s equation for the solute concentration is derived via a matched-asymptotics expansion in the slenderness, $\epsilon$, on the boundary integral solution of Laplace’s equation. This method systematically reduces the problem from solving an implicit indirect surface integral equation over the filament to a significantly simpler explicit calculation of a line integral along the filament centreline. This non-local slender theory for chemically active filaments is valid for arbitrary non-self-intersecting three-dimensional centrelines, and can deal with both varying thickness and non-axisymmetric chemical patterning.
In this work, we use a local version of SPT, consistent with our use of a local slender body theory in the hydrodynamics discussed in Section 2.2. For this, the surface solute concentration is expanded, as in previous work, as
$c(s,\theta)={c}^{(0)}(s,\theta)+\epsilon{c}^{(1)}(s,\theta)+O(\epsilon^{2}).$ (4)
However, for this local theory, we neglect the non-local line integral contributions in each asymptotically derived term, since they are $O(1/\log\epsilon)$ smaller than the local terms. The resulting slip velocity over the surface of the swimmer can then be determined as a surface gradient of the concentration
$\frac{\bm{v}_{\mathrm{slip}}(s,\theta)}{\mathcal{M}(s,\theta)}=$ $\underbrace{\frac{\hat{\bm{e}}_{\theta}(s,\theta)}{\epsilon\rho(s)}\frac{\partial{c}^{(0)}}{\partial\theta}}_{O(1/\epsilon)}+\underbrace{\left[\frac{\hat{\bm{e}}_{\theta}(s,\theta)}{\rho}\frac{\partial{c}^{(1)}}{\partial\theta}+\hat{\bm{t}}(s)\frac{\partial{c}^{(0)}}{\partial s}\right]}_{O(1)}+O(\epsilon),$ (5)
@@
An important caveat here is that reassembling the surface solute concentration from local SPT using eq. 4 is no longer asymptotically accurate, as we neglect some intermediate terms by removing the line integral contributions. However, the resulting dynamics only depend on the slip velocity; crucially, due to the magnification of $\theta$-gradients of ${c}^{(1)}$, we do calculate the leading order slip velocity, since we keep the terms of $O(\log\epsilon)$ and ignore terms of $O(1)$ which are smaller. A comparison of local and non-local SPT is given in Appendix A, showing that local SPT well-approximates the slip velocity.
@@
We focus on the case of an axisymmetric activity, $\mathcal{A}=\mathcal{A}(s)$, where the leading-order and next-order concentrations of local SPT are found to be
@@
where $\kappa$ is the Frenet-Serret curvature and $\Theta$ tracks the azimuthal angle relative to the cumulative torsion (the Frenet-Serret normal coincides with $\Theta=0$).
We note that the leading order surface concentration is uniform over any cross-section, ${c}^{(0)}={c}^{(0)}(s)$, and so the leading-order slip velocity is $O(1)$. The slip velocity calculation can then be further simplified by decomposing $\hat{\bm{e}}_{\theta}$ into Frenet-Serret normal and binormal vectors and considering the cross-sectionally averaged slip-velocity

### v2

@@ Abstract
Active deformable filaments exhibit a large range of qualitatively different three-dimensional dynamics, depending on their flexibility, the strength and nature of the active forcing, and the surrounding environment. We investigate the dynamic behaviour of elastic, chemically propelled phoretic filaments, combining two existing models; a local version of slender phoretic theory, which determines the resulting slip flows for chemically propelled filaments with a given shape and chemical patterning, is paired with a computationally efficient method for capturing the elastohydrodynamics of a deformable filament in viscous flow to study the chemoelastohydrodynamics of filaments. As the activity increases, or equivalently the filament stiffness decreases, these filaments undergo buckling instabilities that alter their behaviour from rigid rods. We follow their behaviour well beyond the buckling threshold to find a rich array of dynamics. Through two illustrative examples, we conduct initial-value simulations that show that, as the stiffness of the filament is decreased, the dynamic behaviour moves from rigid motion to planar buckling, through an out-of-plane transition, eventually reaching diffusive-like behaviours for very deformable filaments.
keywords:
@@ 1 Introduction
Active colloids have a prominent position as self-propelling agents that enable the study of multi-agent systems, called ‘active matter’ (Ramaswamy, 2010). These canonical systems of active matter serve as micro-engineered analogues of crowds, flocks of birds or bacterial swarms. A common propulsion mechanism is phoresis, where particles move in gradients of a surrounding field (Anderson, 1989). This occurs due to short-range interactions between the particle’s surface and the surrounding field that generate an effective slip flow close to the particle’s surface, such as solute concentration (Golestanian et al., 2005), temperature (Bickel et al., 2013) or electric field (Nourhani et al., 2015). If the particle also creates gradients in the surrounding propulsive field, then it can self-propel; this self-propulsion is called autophoresis (Paxton et al., 2004).
One class of autophoretic particles are chemically active (diffusiophoretic) colloids, which are microscale particles that swim by converting a solute ‘fuel’ in their environment into propulsive slip flows (Golestanian et al., 2007). The canonical chemically active colloid is the ‘Janus’ particle, which is half active and half inert; typically this takes the form of a rigid sphere, rod, or disk, partially coated in a catalyst, such as platinum, in a surrounding solute such as hydrogen peroxide solution (Howse et al., 2007; Ebbens & Howse, 2011). Differential reaction rates at the particle’s surface create surface gradients in the solute concentration; the resulting short-range solute-surface interactions generate pressure gradients in a thin layer that drives slip flows close to the surface that propel the swimmer.
@@
Microscale flexible filaments in viscous flows are well-studied, due in part to their prevalence in many biological and artificial systems (Lauga & Powers, 2009; Du Roure et al., 2019). In biology, carpets of slender filaments, known as cilia, actively pump fluid over cycles driven by internal active forces (Ishikawa, 2024); sperm cells swim due to active waves sent down their long and thin tail-like appendages (Gaffney et al., 2011); and bacteria rotate slender, helical flagella to propel themselves (Lauga, 2016). Meanwhile, flexible fibres are present in many industrial and household processes, such as textile fibres in clothing (Duprat, 2022), wood pulp fibres in the production of paper (Lundell et al., 2011), and microplastics that can pollute the environment (DiBenedetto, 2025).
The dynamics of these flexible filaments in viscous flows are difficult to simulate since there is intrinsic feedback between the filament shape, the distribution of forces and the surrounding fluid. Simulations are often computationally expensive because the underlying equations are notoriously numerically stiff, and there are often also accompanying constraints, such as inextensiblity, that must be satisfied throughout the evolution of the system (Du Roure et al., 2019). A range of methods have been used to model and simulate the hydrodynamics of slender flexible filaments. Local slender body theories, such as resistive force theory (Hancock, 1953; Gray & Hancock, 1955; Cox, 1970; Batchelor, 1970; Lighthill, 1976), have been applied to generate fast, efficient simulations of elastic filaments (Moreau et al., 2018; Walker et al., 2020), while non-local slender body theories (Keller & Rubinow, 1976; Johnson, 1977, 1980; Koens & Lauga, 2018) provide greater accuracy through more intensive simulations (Tornberg & Shelley, 2004). Alternative approaches involve using singularity methods such as regularised Stokeslets (Cortez, 2001; Cortez et al., 2005; Hall-McNair et al., 2019), immersed boundary methods (Pozrikidis, 1992; Peskin, 2002) and bead models (Delmotte et al., 2015, and references therein). These models have been developed further to investigate the interaction of flexible fibres and the behaviour of passive and active suspensions (Tornberg & Shelley, 2004; Schoeller et al., 2021).
@@
Active fibres have also been found to undergo similar dynamic transitions depending on the strength of activity relative to the fibre stiffness, with a particular focus on clamped filaments as models of ciliary beating or actin filaments. Bead-spring models of a chain of elastically linked beads, such as with stresslets distributed along its axis (Laskar et al., 2013) or when forced at one end (Laskar & Adhikari, 2017), demonstrated whirling, corkscrewing and planar beating dynamics; similar linear, helical and planar modes were found when attached to a passive cargo (Manna et al., 2017). This bead-spring approach has recently been extended to finer resolution to investigate active poroelastic filaments under ‘follower’ forces that act tangentially to the centreline at the free end of the filament (Altunkeyik et al., 2025), which showed rich transitions between dynamic behaviours, including for chemically active poroelastic filaments, and exhibited regimes of periodic and diffusive-like motion.
@@
An alternative approach to studying active elastohydrodynamics is to combine resistive force theory with the mechanics of elastic filaments. When the motion is confined to lie in the plane, filaments deforming due to follower forces have been shown to undergo Hopf bifurcations, seen as a flapping motion for clamped filaments (De Canio et al., 2017), while a two-dimensional linear stability analysis revealed transitions to rotation or beating that depends on mechanical boundary conditions, applied load, and noise effects (Fily et al., 2020). Further dynamical regimes have been explored using this elastohydrodynamic theory when applying localized tangential forces (Man & Kanso, 2019) to extract motions such as straight swimming, circling, spiralling and undulating, and also with clamped filaments in confined domains (Stein et al., 2021), revealing transitions of filaments with distributed tangential forces from stable to oscillatory and streaming modes. Recent theoretical contributions by Clarke et al. (2024) and Schnitzer (2025) further investigate the full three-dimensional motion of these follower-force models, identifying beating and whirling onset via bifurcation theory and weakly nonlinear analysis, enriching the fundamental understanding of these dynamics. Interestingly, clamped filaments that are free to deform in three-dimensions display a non-planar ‘whirling’ motion following the onset of instability, and only reach a planar beating state when the applied force is much stronger. In a related system, Lough et al. (2023) modelled the swarmer cell state of Proteus mirabilis as an active Kirchoff rod, and found different swimming modes that may be planar or twisted into three-dimensions, depending on the relative material stiffnesses. Although many of these models abstract away the details provided by the surface chemistry of phoretic propulsion, for example by using follower forces, we expect that they retain essential features of force-induced shape transformations and the resulting complex locomotion, and highlight the complex interplay of elasticity, active forcing, and boundary conditions in shaping filament dynamics.
@@
In this paper, we combine recent advances in the efficient simulation of elastohydrodynamic filaments (Walker et al., 2020) with the asymptotically-derived framework for calculating slip flows in chemically active, diffusiophoretic filaments (Katsamba et al., 2020, 2022, 2024) to investigate the fundamental dynamic behaviours of individual chemically propelled elastic filaments. We consider examples of canonical surface chemical patterning to study the transition between qualitatively different dynamic regimes as the material stiffness is varied. Section 2 outlines our theoretical method and numerical implementation for modelling chemically active filaments, using a local slender body theory to calculate the resulting slip flows from chemical reactions on the filament’s surface, which are applied as a boundary condition in the numerical simulations of the elastohydrodynamics. We present our results in section 3, investigating typical behaviours as the material stiffness (or equivalently the activity) is varied, via two key examples: a symmetrically-patterned filament that buckles under its own induced slip flow, and an asymmetrically-patterned filament whose swimming trajectory is altered by its deformation. We discuss and analyse the outcomes further in section 4.
@@ 2 Theory of chemoelastohydrodynamic filaments
We consider an inextensible slender elastic filament in an otherwise quiescent, infinite viscous fluid, as shown in fig. 1. Within the fluid is a chemical solute that diffuses rapidly. The solute acts as a fuel for the filament’s self-propulsion: a chemical reaction of the solute dispersed in the surrounding fluid is catalysed by the catalytic coating on the surface of the filament, and short-range interactions between solute molecules and the filament’s surface generate propulsive slip flows from solute concentration gradients within a thin layer by the filament’s surface. The resulting bulk flow can propel and deform the filament, altering the filament conformation and changing the solute distribution; there is a complex feedback between shape, chemical patterning and swimming trajectory that we wish to explore.
Here, we outline the main theoretical components used to capture the evolution of slender chemically propelled elastic filaments. Our work combines two pre-existing methods: slender phoretic theory (Katsamba et al., 2020, 2022, 2024; Katsamba & Montenegro-Johnson, 2024) that asymptotically calculates the leading order slip flow around slender chemically active filaments when given their chemical patterning and shape, and an efficient method for simulating elastohydrodynamics (Walker et al., 2020). Together, these allow us to probe the dynamic behaviour of chemically propelled elastic filaments, which we refer to as chemoelastohydrodynamics.
2.1 The slender filament geometry
The geometry of a slender filament with a circular cross-section is captured by its centreline position, $\bm{X}(s,t)$, and cross-sectional radius $\rho(s)$, each parametrised by the arclength $s$ such that $|\partial\bm{X}/\partial s|=1$. The filament centreline has a total length $2l$ and the maximum cross-sectional radius is $r_{f}$, and we assume the filament is unshearable, so that cross-sections remains perpendicular to the centreline. We define a slenderness parameter by the ratio of the maximum diameter to the total arclength, $\epsilon=r_{f}/l$, which can be considered an aspect ratio, and we assume that our filaments are long and thin so that $\epsilon\ll 1$.
The surface of the filament is patterned with two important surface chemical properties: the activity, $\mathcal{A}(\bm{x})$, which prescribes a surface solute flux (either generating or depleting solute), and the mobility, $\mathcal{M}(\bm{x})$, that quantifies how chemical gradients drive slip flows. Theoretically, we can set these independently and they may vary over the surface, but we assume the distribution remains fixed in the material frame.
The system is non-dimensionalised: spatial coordinates are rescaled by the half-length of the filament $l$, while the cross-sectional radius is rescaled by the maximum radius $r_{f}$. Taking typical (maximum) values for the activity and mobility as $A$ and $M$, respectively, we also use a concentration scale $Ar_{f}/D$ and a velocity scale $AMr_{f}/Dl$ to rescale our equations, where $D$ is the diffusivity of the solute. All equations are henceforth given in dimensionless form.
@@ 2.2 Local slender phoretic theory
The slip flow generated by the chemical activity is calculated using slender phoretic theory. The full details of this asymptotically-accurate theory are given in Katsamba et al. (2020, 2022, 2024); Katsamba & Montenegro-Johnson (2024); here we outline the key points.
We consider the zero Péclet number limit for the solute dynamics, where diffusion dominates advection, so that the solute concentration, $c$, in the surrounding fluid is governed by Laplace’s equation,
@@
where we consider the disturbance from a background concentration, so that $c\to 0$ far from the surface.
The activity provides a normal flux boundary condition at the filament surface via zeroth order reaction kinetics, and any resulting surface concentration gradients drive the slip flow proportional to the mobility. These are boundary conditions on the surface of the filament
@@
where $\bm{n}_{f}$ is the outward normal to the solid surface. Importantly, we frame our results in terms of the arclength $s$ and local polar coordinates in the cross-section $(r,\theta)$. Surface properties are therefore defined by the arclength $s\in[-1,1]$ along the centreline and angle $\theta\in[-\pi,\pi]$ around the cross-section, i.e. $\mathcal{A}=\mathcal{A}(s,\theta)$ and $\mathcal{M}=\mathcal{M}(s,\theta)$, with the surface at $r=\rho(s)\in[0,1]$. Note that we take $\theta$ as the angle relative to a fixed position in each cross-section in the material frame. These local coordinates are shown in the inset to fig. 1.
In the framework of slender phoretic theory (SPT) (Katsamba et al., 2020, 2022, 2024; Katsamba & Montenegro-Johnson, 2024), the solution of Laplace’s equation for the solute concentration is derived via a matched-asymptotics expansion in the slenderness, $\epsilon$, on the boundary integral solution of Laplace’s equation. This method systematically reduces the problem from solving an implicit indirect surface integral equation over the filament to a significantly simpler explicit calculation of a line integral along the filament centreline. This non-local slender theory for chemically active filaments is valid for arbitrary non-self-intersecting three-dimensional centrelines, and can deal with both varying thickness and non-axisymmetric chemical patterning. The solution for the surface solute concentration for an axisymmetric activity is given in appendix A.
In this work, we use a local version of SPT, consistent with our use of a local slender body theory in the hydrodynamics discussed in Section 2.3. For this, the surface solute concentration is expanded as
$c(s,\theta)={c}^{(0)}(s,\theta;\epsilon)+\epsilon{c}^{(1)}(s,\theta;\epsilon)+o(\epsilon),$ (4)
where the terms ${c}^{(0)},~{c}^{(1)}$ are allowed to depend weakly (i.e. logarithmically) upon $\epsilon$, and we calculate only the leading order component of each of these terms.This corresponds to neglecting the non-local line integral contributions in each asymptotically derived term of the full slender phoretic theory (quoted in appendix A), since they are $O(1/\log\epsilon)$ smaller than the local terms. The resulting slip velocity over the surface of the swimmer can then be determined as a surface gradient of the concentration
$\frac{\bm{v}_{\mathrm{slip}}(s,\theta)}{\mathcal{M}(s,\theta)}=$ $\underbrace{\frac{\hat{\bm{e}}_{\theta}(s,\theta)}{\epsilon\rho(s)}\frac{\partial{c}^{(0)}}{\partial\theta}}_{O(1/\epsilon)}+\underbrace{\left[\frac{\hat{\bm{e}}_{\theta}(s,\theta)}{\rho}\frac{\partial{c}^{(1)}}{\partial\theta}+\hat{\bm{t}}(s)\frac{\partial{c}^{(0)}}{\partial s}\right]}_{O(1)}+o(1),$ (5)
@@
Local slender phoretic theory calculates the leading order slip velocity, which is required to determine the leading order behaviour of the resulting dynamics. However, it is important to note that reassembling the surface solute concentration from local SPT using eq. 4 is no longer asymptotically accurate, as we neglect some intermediate terms by removing the line integral contributions. Crucially, due to the magnification of $\theta$-gradients of ${c}^{(1)}$, the leading order slip velocity is accurately calculated, since we retain terms of $O(\log\epsilon)$ and ignore terms of $O(1)$ which are smaller. A comparison of local and non-local SPT is given in Appendix B, showing that local SPT well-approximates the slip velocity.
@@
We focus on the case of an axisymmetric activity, $\mathcal{A}=\mathcal{A}(s)$. For this form of chemical patterning, in both local and non-local slender phoretic theories, the leading order surface solute concentration is also axisymmetric and, hence, independent of $\theta$: ${c}^{(0)}(s,\theta)={c}^{(0)}(s)$ (Katsamba et al., 2020). In fact, the leading order and next order concentrations of local SPT are found to be
@@
where $\kappa$ is the Frenet-Serret curvature and $\Theta$ is the azimuthal angle relative to the Frenet-Serret basis, tracking the azimuthal angle relative to the cumulative torsion (note that $\Theta$ is not fixed relative to the material frame and the Frenet-Serret normal coincides with $\Theta=0$). We note that the leading order surface concentration is uniform over any cross-section, ${c}^{(0)}={c}^{(0)}(s)$, and so the leading order slip velocity is $O(1)$.
To determine the leading order kinematics of the filament, it is only necessary to calculate the cross-sectional average of the slip velocity, since all higher order modes in a Fourier series in $\theta$ have smaller effect in the motion of the slender body (see, for example, Katsamba et al., 2020, §2.9). The slip velocity calculation can then be further simplified by decomposing $\hat{\bm{e}}_{\theta}$ into Frenet-Serret normal and binormal vectors, so that we find the cross-sectional average slip velocity to be